OpenStudy (anonymous):

let f(x)=x^2-3x+1 and g(x)=3x-1. find f(g(x)) I got: 9x^2-9x-3

OpenStudy (anonymous):

f(g(x))=(3x-1)^2-3(3x-1)+1 =(3x-1)(3x-1)-9x+3+1 =9x^2-6x+1-9x+4 =9x^2-15x+5

OpenStudy (anonymous):

dangit. i was close.:P